ISOMER-SPECIFIC INFRARED SPECTROSCOPY OF IONIC COMPLEXES

\underline{JOSEPH R. ROSCIOLI}, BEN ELLIOTT, MARK A. JOHNSON, Sterling Chemistry Laboratory, Yale University, PO Box 208107, New Haven, CT 06520.

We report a new technique which allows for isomer-specific infrared spectroscopy of gas-phase ionic species. In this twolaser, tandem time-of-flight method, we monitor the ground-state population of a specific isomer by tuning an infrared laser to one of its vibrational transitions and measuring the resulting predissociation yield. This population is depleted when another, preceding infrared laser scans through the vibrational resonances of that isomer, thus isolating its infrared spectrum. We demonstrate the generality of this predissociation-dip method by applying it to several ionic system, including anionic Cl^{−}^{.}HDO^{.}Ar and cationic Aniline^{.}H^{+}^{.}Ar. In the hydrated chloride study, we separate the infrared spectrum of Cl^{−}^{.}HOD (where the H is shared between the Cl^{−} and O) from that of Cl^{−}^{.}DOH, and empirically assign a band near 3000 cm^{−}^{1} to the Cl^{−}^{.}HOD isomer. We will also discuss the utility of this technique in the context of recent experimental studies of large proton-bound dimers.